if tom and lynn were working on a garden together. and tom was able to prepare a whole garden in about 8 hours and lynn was able to do it in 12 hours. how long would it take for both tom and lynn working together to prepare a whole garden?
Answer by richwmiller(17219)

1/a+1/b=1/x
1/8+1/12=1/x
12/96+8/96=1/x
20/96=1/x
x=96/20
x=4.8 hours working together
check
x/a+x/b=1
x/8+x/12=1
4.8/8+4.8/12=1
0.6+0.4=1
1.0=1
ok
codewt